OpenStudy (anonymous):

1) What is the value of 7p4? A. 28 B. 120 C. 210 D. 840 2) Out of six students in a class, three will be chosen for student government positions. What is the number of permutations of those six students to hold the three different positions? A. 3 B. 18 C. 81 D. 120 3) A restaurant offers 7 different appetizers and 11 different entrees. What is the number of possible unique combinations of appetizer and entree that can be ordered? A. 18 B. 77 C. 110 D. 770

OpenStudy (anonymous):

Option C. 210

OpenStudy (anonymous):

How did you get that??

OpenStudy (anonymous):

1) 7*6*5*4=840

OpenStudy (anonymous):

Okay what about 2 and 3?

OpenStudy (anonymous):

2)6p3=6*5*4=120

OpenStudy (anonymous):

#1- D (So sorry about it) , #2-D ... well 7p4 = 7!/(7-4)! and the next shall be 6c3 *3! where 6c3= 6!/((6-3)!*3!)

OpenStudy (anonymous):

Thankss and what about 3?

OpenStudy (anonymous):

I don't think the options are correct, mate. The answer should be 7^11. But, ask this question in chat to see if anybody else can help you out. Maybe, I'm wrong.

OpenStudy (anonymous):

ok thanks anyways.